# Dose-aware Diffusion Model for 3D Low-count PET: Multi-institutional Validation with Reader Study and Real Low-dose Data

### v2

@@ III-D Comparison with other baseline denoising methods
Input 44.667 / 0.682 / 0.788 49.260 / 0.404 / 0.895 53.575 / 0.249 / 0.957 56.085 / 0.189 / 0.977 59.346 / 0.131 / 0.991 62.253 / 0.093 / 0.996
UNN 52.034 / 0.293 / 0.953 54.044 / 0.235 0.973 55.838 / 0.194 / 0.982 56.963 / 0.173 / 0.987 58.570 / 0.146 / 0.991 59.885 / 0.126 / 0.994
DDIM 42.550 / 0.852 / 0.904 42.718 / 0.836 / 0.926 42.846 / 0.824 / 0.938 42.942 / 0.815 / 0.944 43.065 / 0.804 / 0.952 43.145 / 0.797 / 0.957
DiffusionMBIR 42.590 / 0.848 / 0.913 42.747 / 0.833 / 0.934 42.868 / 0.822 / 0.944 42.960 / 0.814 / 0.950 43.079 / 0.803 / 0.957 43.156 / 0.796 / 0.961
TPDM 42.691 / 0.839 / 0.906 42.810 / 0.827 / 0.929 42.939 / 0.815 / 0.940 43.051 / 0.805 / 0.946 43.202 / 0.791 / 0.954 43.291 / 0.783 / 0.959
DDPET-3D (proposed) 52.899 / 0.267 / 0.965 54.937 / 0.215 / 0.977 57.119 / 0.171 / 0.985 58.551 / 0.148 / 0.989 60.916 / 0.117 / 0.993 63.804 / 0.088 / 0.996
The proposed DDPET-3D was compared with UNN [16], DiffusionMBIR [27], and TPDM [28]. As described previously, UNN was chosen because it was also proposed to achieve noise-aware denoising and was one of the winning methods in 2022 Ultra Low-dose PET Imaging Challenge. DiffusionMBIR and TPDM were selected because they were recent works proposed to address the diffusion 3D inconsistencies problem in medical imaging. We also compared with the standard DDIM (Denoising Diffusion Implicit Models) sampling [31]. Detailed implementations of different methods will be provided later in this paper (see Methods).
As presented in Fig. 9, standard DDIM sampling produced images with severe inconsistencies between slices in the sagittal view. Both DiffusionMBIR and TPMD improved the results, but the problem still exists. Proposed DDPET-3D method produced noticeably more consistent reconstructions. Moreover, even though other diffusion models produced visually realistic PET images, they failed to accurately reconstruct the structure of different organs. For example, all other diffusion models produced images with distorted myocardium as shown in the transverse slice in Fig. 9. The proposed DDPET-3D method also recovered some fine details in the images, such as the hot spot pointed by the blue arrows in Fig. 9, which is either over-smoothed (in UNN) or barely visible (in other diffusion methods).
Another major issue of the diffusion model is inaccurate image quantification. Note that all the presented images were already normalized by the total injected activities. As shown in Fig. 9, the tracer activities in certain organs are completely wrong in other existing diffusion models. For example, the activities in the brain are noticeably lower in DDIM, DiffusionMBIR, and TPDM results; the myocardium is distorted in competing baseline methods. Such differences may affect certain diagnostic tasks such as lesion detection [5]. Even though UNN produced over-smoothed reconstructions, it does not alter the overall tracer activities in different organs. Using UNN output as denoised prior (see details in Methods), the proposed DDPET-3D maintained overall image quantification. Also, using UNN as a denoised prior allows the proposed method to recover some subtle features that are almost invisible in the low-count input (blue arrows in Fig. 9).
Images were quantitatively evaluated using SSIM (Structural Similarity Index), PSNR (Peak signal-to-noise ratio), and NRMSE (Normalized Root-mean-square error). To facilitate the testing process, we tested all the comparison methods using 20 patients from the entire testing dataset. There were 6 different low-count levels of all the patients, resulting in a total of 120 testing studies. Quantitative results are presented in Table II.
As presented in Table II, the proposed method outperformed other methods in all the 6 low-count levels. When the input count-level increases, the performance of all the methods gradually improves. Both the proposed method and UNN can achieve noise-aware denoising. However, at higher count-levels (25% and 50%), UNN results were even worse than the input. In contrast, the proposed method produced optimal results at all count-levels. Also, due to inaccurate image quantification, other diffusion models produced images with even worse quantitative results compared to low-count inputs at higher count-levels.
As presented in Table II, UNN is the second-best method. To comprehensively evaluate the proposed method, We compared the proposed method with UNN for all the 1,137 testing patients across three different scanners. As presented in Table I, DDPET-3D consistently outperformed UNN on different low-count levels.
III-E Albation studies
@@
TABLE III: Quantitative assessment for different ablation studies. The best results among different low-count levels are marked in red. The second best results are marked in blue. Black voxels were removed for calculations. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
@@
DDPET-3D (9 slices) 52.563 / 0.277 / 0.964 54.666 / 0.215 / 0.977 56.194 / 0.162 / 0.984 57.688 / 0.137 / 0.988 60.011 / 0.106 / 0.993 62.267 / 0.082 / 0.995
DDPET-3D (21 slices) 52.502 / 0.278 / 0.960 54.509 / 0.224 / 0.974 56.574 / 0.181 / 0.981 57.949 / 0.157 / 0.985 60.137 / 0.125 / 0.990 62.338 / 0.100 / 0.994
DDPET-3D (41 slices) 51.866 / 0.262 / 0.964 53.926 / 0.208 / 0.977 56.147 / 0.162 / 0.985 57.618 / 0.138 / 0.989 59.948 / 0.107 / 0.993 62.438 / 0.081 / 0.995
DDPET-3D (no prior) 41.497 / 0.961 / 0.868 41.638 / 0.945 / 0.898 41.704 / 0.938 / 0.913 41.715 / 0.937 / 0.920 41.747 / 0.934 / 0.927 41.767 / 0.932 / 0.932
DDPET-3D (no fix $\epsilon$) 52.991 / 0.264 / 0.961 55.049 / 0.212 / 0.973 57.225 / 0.170 / 0.980 58.695 / 0.146 / 0.983 61.026 / 0.116 / 0.988 63.916 / 0.087 / 0.991
DDPET-3D (no dose) 52.817 / 0.269 / 0.960 54.822 / 0.217 / 0.972 56.903 / 0.175 / 0.979 58.259 / 0.152 / 0.983 60.360 / 0.123 / 0.987 62.871 / 0.096 / 0.990
DDPET-3D (single $\epsilon$) 52.289 / 0.285 / 0.959 54.261 / 0.231 / 0.974 56.324 / 0.185 / 0.983 57.701 / 0.161 / 0.987 59.917 / 0.128 / 0.992 62.548 / 0.098 / 0.995
DDPET-3D (proposed) 52.899 / 0.267 / 0.965 54.937 / 0.215 / 0.977 57.119 / 0.171 / 0.985 58.551 / 0.148 / 0.989 60.916 / 0.117 / 0.993 63.804 / 0.088 / 0.996
TABLE IV: Quantitative assessment for different methods. The measurements were obtained by averaging the values on the testing human studies. The proposed method consistently produced promising denoised results regardless of input count levels. $p<0.01$ was observed in all groups when comparing the proposed method with all other methods (except for 50% SSIM values). The best results among different low-count levels are marked in red. Black voxels were removed for calculations. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric.
@@
Input 44.667 / 0.682 / 0.788 49.260 / 0.404 / 0.895 53.575 / 0.249 / 0.957 56.085 / 0.189 / 0.977 59.346 / 0.131 / 0.991 62.253 / 0.093 / 0.996
UNN 52.034 / 0.293 / 0.953 54.044 / 0.235 / 0.973 55.838 / 0.194 / 0.982 56.963 / 0.173 / 0.987 58.570 / 0.146 / 0.991 59.885 / 0.126 / 0.994
DDIM50 42.550 / 0.852 / 0.904 42.718 / 0.836 / 0.926 42.846 / 0.824 / 0.938 42.942 / 0.815 / 0.944 43.065 / 0.804 / 0.952 43.145 / 0.797 / 0.957
DDIM50+2.5D 43.636 / 0.754 / 0.910 43.851 / 0.736 / 0.924 43.856 / 0.735 / 0.930 43.829 / 0.738 / 0.932 43.805 / 0.740 / 0.936 43.820 / 0.738 / 0.938
DiffusionMBIR 42.590 / 0.848 / 0.913 42.747 / 0.833 / 0.934 42.868 / 0.822 / 0.944 42.960 / 0.814 / 0.950 43.079 / 0.803 / 0.957 43.156 / 0.796 / 0.961
DiffusionMBIR+2.5D 42.615 / 0.846 / 0.913 42.834 / 0.825 / 0.935 42.892 / 0.819 / 0.945 42.911 / 0.818 / 0.948 42.931 / 0.816 / 0.952 42.950 / 0.814 / 0.955
TPDM 42.691 / 0.839 / 0.906 42.810 / 0.827 / 0.929 42.939 / 0.815 / 0.940 43.051 / 0.805 / 0.946 43.202 / 0.791 / 0.954 43.291 / 0.783 / 0.959
TPDM+2.5D 42.334 / 0.873 / 0.887 42.458 / 0.861 / 0.912 42.483 / 0.859 / 0.923 42.499 / 0.857 / 0.927 42.528 / 0.854 / 0.932 42.536 / 0.853 / 0.935
DDPET-3D (proposed) 52.899 / 0.267 / 0.965 54.937 / 0.215 / 0.977 57.119 / 0.171 / 0.985 58.551 / 0.148 / 0.989 60.916 / 0.117 / 0.993 63.804 / 0.088 / 0.996
@@
Figure 14: Without the denoised prior, proposed DDPET-3D still produced more consistent reconstructions compared with other methods. Blue arrows point to regions with noticeable inconsistencies between slices in competing methods.
@@
Impact of Number of Conditional Slices: We evaluated the DDPET-3D with different numbers of neighboring conditional slices. Three variants of DDPET-3D networks were trained using 9, 21, and 41 neighboring slices as conditional information. These networks are denoted as ”DDPET-3D ($n$ slice)”, where $n$ is the number of conditioned neighboring slices. Results showed that $n=31$ performed the best in most quantitative metrics. Therefore, $n=31$ was used in this paper. Quantitative results are presented in Table III. As presented in Fig. 10, using more conditioned slices helped the network to recover some subtle details in the images (blue arrows in Fig. 10).
@@
For comparison, we also extended the comparison diffusion models (DiffusionMBIR, TPDM, and DDIM), and re-trained them using neighboring slices as conditional information (denoted as DiffusionMBIR+2.5D, TPDM+2.5D, and DDIM+2.5D). Sample reconstruction results are presented in Fig. 11. Simply adding neighboring slices as conditional information does not necessarily lead to a better performance in other baseline models. Using the same testing patients in Table II, corresponding quantitative evaluations are presented in Table IV.
@@
Impact of Denoised Prior: We tested the DDPET-3D without the denoised prior during the sampling process to demonstrate the improvement in image quantification using the proposed denoised prior. This network is denoted as ”DDPET-3D (no prior)” in Table III. As presented in Fig. 12, we can see that diffusion models produced images with inaccurate tracer activities in different organs without the proposed denoised prior (especially in the brain and liver). Note that the images were already normalized by the total injected activities of the entire 3D volume. Also, without the denoised prior, many details in the images were not able to be recovered, as presented in the brain slice in Fig. 12. In addition, we implemented the denoised prior in other competing methods for comparison, so that all the methods have the same starting point in the diffusion process. As shown in Fig. 13, DDPET-3D still outperformed other competing methods, demonstrating the effectiveness of other proposed techniques in DDPET-3D. With the denoised prior, DiffusionMBIR and TPDM produced images with better PSNR and RMSE measurements than the networks without it, but inconsistencies between slices still exist. Lastly, as shown in Fig. 14, without the denoised prior, DDPET-3D still produced more consistent reconstructions compared with other methods. Note that the denoised prior is one of the key contributions proposed in this work.
@@
Impact of Fixing Noise Variables: We analyzed the DDPET-3D without and with fixing the 2 noise variables to demonstrate its effectiveness in producing consistent 3D reconstructions. Specifically, $\epsilon_{0}^{a}$ and $\epsilon_{0}^{b}$ are sampled from ${\mathcal{N}}(0,\textbf{I})$ for every slice in the 3D image volume, instead of fixing them for all slices. This network is denoted as ”DDPET-3D (no fix $\epsilon$)” in Table III. Although the differences in quantitative evaluations without and with fixing noise variables are small, and not fixing noise even led to better quantitative measurements in certain cases, we noticed significant improvements in visual quality, as presented in Fig. 15. Not fixing noise variables produced images with inconsistent slices and unclear organ boundaries, which are unfavorable in clinical settings.
@@
Impact of Dose Embedding: We tested the DDPET-3D without dose embedding, which is denoted as ”DDPET-3D (no dose)” in Table III. Experimental results showed that DDPET-3D with dose embedding produced images with better quantitative results.
IV Discussion and Conclusion

### v3

@@ IV-D Generalizability test in a different hospital
DDPET-3D (Trained on Ruijin) 54.590* / 0.264* / 0.965* 56.267* / 0.222* / 0.975* 58.215* / 0.150* / 0.985 59.470 / 0.143* / 0.989 61.259* / 0.115* / 0.993 63.650 / 0.087 / 0.996
DDPET-3D (Trained on Bern) 55.909 / 0.191 / 0.980 57.630 / 0.154 / 0.987 59.370 / 0.127 / 0.991 60.631 / 0.113 / 0.993 62.229 / 0.093 / 0.995 64.237 / 0.074 / 0.996
[…]
DDPET-3D (Trained on Ruijin) 49.673* / 0.243* / 0.945* 51.195 / 0.206 / 0.959* 53.453* / 0.157* / 0.969*
DDPET-3D (Trained on Yale) 49.950 / 0.262 / 0.953 51.383 / 0.199 / 0.966 53.056 / 0.165 / 0.976
[…]
DDPET-3D (Trained on Ruijin) 50.134 / 0.179 / 0.962 51.964 / 0.145 / 0.973 53.348 / 0.123 / 0.979 54.772 / 0.102 / 0.987 56.607 / 0.085 / 0.991
DDPET-3D (Trained on Davis)
To evaluate the cross-center generalizability of the proposed DDPET-3D, we applied the model trained using the uExplorer data acquired at Shanghai Ruijin Hospital in China to patient studies acquired at the other three hospitals.
IV-D1 Same Scanner, Different Hospitals
We first examine the generalizability of the data from UC Davis Medical Center in the USA. Although both hospitals used the same scanner model, there are numerous factors that could lead to generalization challenges.
In addition to different patient characteristics in different countries, two hospitals also implement different reconstruction parameters, reconstruction voxel sizes, and acquisition protocols (see details in Methods). Images acquired at different hospitals also have varied image contrast due to differences in scan time and acquisition start time. Shanghai Ruijin Hospital implemented a total of 5 min of scan duration. At UC Davis hospital, they implemented a total of 20 min of scan duration, similar to a typical multi-bed position whole-body scan. Additionally, UC Davis Medical Center also doubled the conventional tracer uptake time from 60 min post-injection to 120 min post-injection for improved tumor to background ratio in oncologic imaging. Shanghai Ruijin Hospital implemented the conventional 60 min tracer post-injection time.
Ten patient studies from the UC Davis Medical Center were included in this experiment. Five low-count levels were generated through listmodel rebinning, including 2.5%, 6.25%, 12.5%, 25%, and 50%. Sample denoised results acquired at the UC Davis Medical Center are presented in Fig. 8. The proposed DDPET-3D showed superior generalizability when directly applied to data acquired at a different hospital. The corresponding quantitative measurements are included in Table II.
Another reader study was conducted to evaluate the overall image quality of the 10 UC Davic patients. Similarly, 25% and 50% low-count and denoised images, as well as 100% full-count images, were included for reader study. As presented in Fig. 9, even though the DDPET-3D denoised images have similar quantitative measurements to low-count images shown in Table I at 25% and 50%, all three readers unanimously gave superior overall rankings to 25% and 50% denoised results than the 100% full-count images. Since the quantitative values shown in Table I were calculated using 100% full-count as reference, they may not correspond to the clinical performance and physicians’ preferences. While the DDPET-3D denoised images appear slightly over-smoothed in Fig. 8 compared to full-count images, probably because the model was not fine-tuned for UC Davis data, the reader study still favored them for their improved noise suppression, demonstrating the superior generalizability of the proposed DDPET-3D. However, further validation against clinical gold standards is needed to fully demonstrate its clinical potential across diverse institutions and imaging protocols.
IV-D2 Different Scanners, Different Hospitals
We also applied the trained DDPET-3D model to other hospitals to assess its generalizability. As shown in Table II, despite slight performance degradations, the trained DDPET-3D model demonstrated strong performance across different hospitals. In some cases, the performance differences between DDPET-3D models trained on local hospital data and those trained on data from Shanghai Ruijin hospital were not statistically significant ($p>0.05$), indicating good generalizability across institutions.
These results suggest that DDPET-3D might be transferable to another medical center without further network fine-tuning, though further investigations are needed.
IV-E Lesion Detectability Evaluation
Figure 10: Lesion detectability results on different low-count levels. Arrows point to the three Monte-Carlo simulated lesions at varying sizes. Denoised images at different low-count levels are presented. The lesion detectability index ($d$) and SUV statistics are included in the figure.
The previous subsections demonstrated the superior denoising performance of DDPET-3D. Another key challenge in PET image denoising is maintaining lesion detectability while removing excessive noise. In this subsection, we attempt to evaluate lesion quantification before and after denoising.
IV-E1 Based on Monte-Carlo simulated lesions
Accurately quantifying lesion detectability in vivo is challenging because there is no known ground truth for lesion uptake. To address this issue and evaluate the lesion detectability of the denoised images generated by DDPET-3D, we leveraged a Monte Carlo (MC)-based tool to simulate the listmode data of spherical lesions of varying sizes of 4mm, 6mm, and 8mm in diameter. The lesions were placed in a straight line in the same coronal image slice in the spleen. The simulated listmode data were inserted into the raw data of a lesion-absent 73-year-old female (BMI: 24.8) scanned with the United Imaging uExplorer total-body PET/CT system. An activity concentration ratio of 2.5:1 was simulated with respect to the background. The images were reconstructed using the UC Davis in-house OSEM algorithm with 4 iterations and 13 subsets without point spread function modeling. The reconstructed voxel size was $1.67\text{mm}\times 1.67\text{mm}\times 2.89\text{mm}$.
As detailed in [44], spheres in air were simulated using the MC-toolkit SimSET  [^2] SimSET Homepage [45] with 500 million 18F decays. SimSET has previously been successfully employed for scatter correction [46] and for image reconstruction algorithms [47] in total-body PET. Patient attenuation, scatter, detector normalization, and decay during the measurement were taken into account. Based on the background activity concentration in the spleen, a subset of the total number of simulated events was chosen to correctly reflect the desired activity concentration ratio of 2.5:1. The simulated events were then merged with the measured data through random shuffling, creating a realistic listmode dataset.
As shown in Fig. 10, three noisy images were reconstructed at different frame lengths: 120s, 300s and 600s, corresponding to approximately 10%, 25%, and 50% of the full count level, respectively, at the UC Davis Medical Center. The lesion detectability index was quantified using a non-prewhitening matched filter [48]. Specifically, the detectability index was calculated as $d=\lambda_{1}-\lambda_{0}$, where $\lambda_{i}$ is the cross-correlation between the lesion-present ($i=1$) and lesion-absent ($i=0$) images, respectively, and a matched filter $w$. $w$ is defined as the difference between the lesion-present and lesion-absent ground-truth images. As demonstrated by the calculated lesion detectability indices $d$ shown in Fig. 10, the proposed DDPET-3D method effectively improved/maintained lesion detectability for this simulated case. Based on the small volumes of interest (VOI) placed on the simulated lesions, the denoised images generated by DDPET-3D exhibited lower $\mathrm{SUV_{max}}$ and higher $\mathrm{SUV_{mean}}$, indicating effective noise suppression and improved lesion contrast.
TABLE III: Absolute percentage difference in lesion-level SUV statistics for different methods across all test patients from four hospitals. Comparisons are made relative to the full-dose reference. The best results among different low-count levels are marked in red. Asterisks (*) denote statistically significant differences ($p<0.05$) between the comparison methods and DDPET-3D.
Lesion SUV Mean / Max Error (%) 1% Count Input 2% Count Input 5% Count Input 10% Count Input 25% Count Input 50% Count Input
UNN 36.0* / 44.1* 26.6* / 36.5* 17.3* / 28.9* 13.1* / 24.8* 8.4* / 19.0* 4.3* / 11.6*
DDIM 62.9* / 58.5* 58.9* / 58.0* 54.8* / 55.8* 53.2* / 55.6* 50.4* / 53.5* 47.1* / 54.0*
DiffusionMBIR 64.8* / 64.4* 60.6* / 63.5* 56.5* / 62.0* 54.4* / 60.6* 51.7* / 59.0* 48.9* / 58.2*
TPDM 65.1* / 64.7* 60.6* / 62.3* 57.0* / 61.3* 57.4* / 60.6* 53.3* / 59.2* 49.5* / 58.0*
Spach Transformer 49.0* / 43.4* 34.5* / 36.1* 21.3* / 27.1* 15.1* / 23.1* 10.3* / 20.1* 6.2* / 15.4*
AMIR 57.9* / 64.9* 42.4* / 56.4* 27.6* / 49.9* 20.3* / 47.2* 14.1* / 45.2* 12.7* / 46.4*
DDPET-3D (proposed) 33.7 / 39.0 21.1 / 28.6 11.4 / 20.4 8.0 / 17.0 5.2 / 13.5 2.9 / 8.7
Lesion SUV Mean / Max Error (%) 1% Count Input 2% Count Input 5% Count Input 10% Count Input 25% Count Input 50% Count Input
UNN 22.0* / 27.1 13.0* / 16.7 6.5 / 13.5 5.3 / 13.2* 4.0 / 10.8 3.5 / 8.1
DDIM 68.7* / 63.0* 65.8* / 62.4* 60.2* / 56.3* 59.4* / 56.1* 55.3* / 54.3* 52.9 / 52.5
DiffusionMBIR 71.9* / 73.4* 68.5* / 69.3* 65.6* / 66.3* 62.6* / 65.3* 61.5* / 64.8* 59.0* / 60.7*
TPDM 69.1* / 70.6* 65.5* / 66.3* 61.2* / 61.9* 57.0* / 58.5* 53.6* / 55.5* 51.3* / 55.2*
Spach Transformer 26.2* / 24.9 15.2* / 17.5 7.7* / 13.9 5.8 / 11.7 4.3 / 9.8 3.3 / 8.1
AMIR 37.4* / 51.2* 27.6* / 46.6* 19.6* / 43.8* 16.0* / 42.5* 13.6* / 41.8* 12.5* / 41.6*
DDPET-3D (proposed) 17.6 / 26.2 11.0 / 16.2 6.4 / 13.5 4.9 / 11.6 4.2 / 11.6 3.2 / 8.1
Lesion SUV Mean / Max Error (%) 5% Count Input 10% Count Input 20% Count Input
UNN 14.4* / 20.5* 9.0 / 12.8 8.5* / 10.6*
DDIM 61.1* / 51.8* 56.8* / 50.6* 54.9* / 48.0*
DiffusionMBIR 61.2* / 59.0* 57.9* / 56.3* 55.9* / 55.6*
TPDM 60.6* / 57.3* 59.3* / 58.0* 57.3* / 55.2*
Spach Transformer 20.3* / 21.8* 12.6* / 13.7* 7.1* / 10.3*
AMIR 26.1* / 41.3* 19.3* / 37.8* 14.4* / 34.8*
DDPET-3D (proposed) 12.0* / 18.0* 8.4 / 12.0 5.3 / 8.4
Lesion SUV Mean / Max Error (%) 2.5% Count Input 6.25% Count Input 12.5% Count Input 25% Count Input 50% Count Input
UNN 32.1* / 32.2* 21.8* / 21.7* 16.7* / 17.1* 11.1* / 11.1 7.4 / 8.1
DDIM 64.9* / 54.5* 59.1* / 56.9* 56.3* / 51.3* 50.6* / 38.8* 47.1* / 38.7*
DiffusionMBIR 67.1* / 65.1* 61.1* / 64.1* 56.8* / 54.5* 55.0* / 56.2* 49.5* / 44.1*
TPDM 68.3* / 70.3* 63.9* / 65.7* 58.0* / 58.7* 58.3* / 54.3* 59.2* / 51.0*
Spach Transformer 35.1* / 33.1* 24.1* / 22.0* 16.8* / 17.2* 12.0* / 14.0 9.7* / 13.4*
AMIR 47.7* / 45.9* 36.6* / 35.4* 27.0* / 29.4* 18.7* / 25.7 11.1* / 24.2
DDPET-3D (proposed) 29.8 / 23.0 17.7 / 13.2 12.2 / 8.8 5.9 / 9.3 4.3 / 7.4
IV-E2 Based on a lesion segmentation network
To achieve a large-scale evaluation on lesion detectability, we employed an nnU-Net [49] model to generate lesion segmentation masks for follow-up lesion SUV analysis. The nnU-Net was first pre-trained using the autoPET dataset [50] and then fine-tuned using the lesion segmentation labels annotated by our board-certified nuclear medicine physicians. Due to limited annotation resources, fine-tuning was performed only on data acquired from Yale New Haven Hospital. Despite being fine-tuned solely on Yale data, the nnU-Net generalized well to patient studies from other hospitals in terms of lesion segmentation. The fine-tuned model was subsequently applied to patient studies from all hospitals to automatically segment potential lesions and generate lesion volumes of interest (VOIs). We then compared lesion-level SUV statistics across different methods against the corresponding full-dose reference data. This large-scale evaluation complements the simulation study and further supports the robustness and clinical applicability of our approach.
The corresponding results on lesion SUV quantification are summarized in Table III. The results demonstrate that the proposed DDPET-3D method achieved consistent and accurate lesion quantification across multiple hospitals, outperforming existing approaches at different low-count levels.
IV-F Ablation studies
@@
TABLE IV: Quantitative assessment for different ablation studies. The reported metrics–PSNR, NRMSE, and SSIM–were computed by averaging values over the entire testing dataset. The best results among different low-count levels are marked in red. The second best results are marked in blue. Black voxels were removed for the calculations. $\uparrow$ and $\downarrow$ indicate that the higher and lower values are better, respectively, for each metric. Asterisks (*) denote statistically significant differences ($p<0.05$) between the comparison methods and the proposed DDPET-3D.
@@
DDPET-3D (9 slices) 52.563* / 0.244* / 0.964 54.666* / 0.193 / 0.977 56.194 / 0.148 / 0.984 57.688 / 0.126 / 0.988 60.011 / 0.097 / 0.993 62.267* / 0.076 / 0.995
DDPET-3D (21 slices) 52.502* / 0.246* / 0.960* 54.509* / 0.198* / 0.974* 56.574* / 0.160* / 0.981* 57.949* / 0.139* / 0.985* 60.137* / 0.111* / 0.990* 62.338* / 0.088* / 0.994*
DDPET-3D (41 slices) 51.866 / 0.238 / 0.964 53.926 / 0.189 / 0.977 56.147 / 0.149 / 0.985 57.618 / 0.127 / 0.989 59.948 / 0.099 / 0.993 62.438* / 0.075 / 0.995
DDPET-3D (no prior) 41.497* / 0.847* / 0.868* 41.638* / 0.833* / 0.898* 41.704* / 0.827* / 0.913* 41.715* / 0.826* / 0.920* 41.747* / 0.823* / 0.927* 41.767* / 0.821* / 0.932*
DDPET-3D (no fix $\epsilon$) 52.991* / 0.233* / 0.961* 55.049* / 0.187* / 0.973* 57.225* / 0.150* / 0.980* 58.695* / 0.129* / 0.983* 61.026* / 0.102 / 0.988* 63.916* / 0.077 / 0.991*
DDPET-3D (no dose) 52.817* / 0.237* / 0.960* 54.822* / 0.192* / 0.972* 56.903* / 0.154* / 0.979* 58.259* / 0.134* / 0.983* 60.360* / 0.108* / 0.987* 62.871* / 0.084* / 0.990*
DDPET-3D (single $\epsilon$) 52.289* / 0.251* / 0.959* 54.261* / 0.203* / 0.974* 56.324* / 0.164* / 0.983* 57.701* / 0.142* / 0.987* 59.917* / 0.113* / 0.992* 62.548* / 0.087* / 0.995*
DDPET-3D (proposed) 52.899 / 0.267 / 0.965 54.937 / 0.215 / 0.977 57.119 / 0.171 / 0.985 58.551 / 0.148 / 0.989 60.916 / 0.103 / 0.993 63.804 / 0.077 / 0.996
TABLE V: Quantitative assessment for different methods. The reported metrics–PSNR, NRMSE, and SSIM–were computed by averaging values over the entire testing dataset. The proposed method consistently produced promising denoised results regardless of input count levels. The best results among different low-count levels are marked in red. Black voxels were removed for calculations. $\uparrow$ and $\downarrow$ indicate that higher and lower values are better, respectively, for each metric. Asterisks (*) denote statistically significant differences ($p<0.05$) between the comparison methods and the proposed DDPET-3D.
@@
Input 44.667* / 0.600* / 0.788* 49.260* / 0.356* / 0.895* 53.575* / 0.219* / 0.957* 56.085* / 0.167* / 0.977* 59.346* / 0.115 / 0.991* 62.253* / 0.082 / 0.996*
UNN 52.034* / 0.258* / 0.953* 54.044* / 0.207* / 0.973* 55.838* / 0.171* / 0.982* 56.963* / 0.152* / 0.987* 58.570* / 0.128* / 0.991* 59.885* / 0.111* / 0.994*
DDIM50 42.550* / 0.751* / 0.904* 42.718* / 0.737* / 0.926* 42.846* / 0.726* / 0.938* 42.942* / 0.718* / 0.944* 43.065* / 0.708* / 0.952* 43.145* / 0.702* / 0.957*
DDIM50+2.5D 43.636* / 0.665* / 0.910* 43.851* / 0.648* / 0.924* 43.856* / 0.648* / 0.930* 43.829* / 0.650* / 0.932* 43.805* / 0.652* / 0.936* 43.820* / 0.650* / 0.938*
DiffusionMBIR 42.590* / 0.748* / 0.913* 42.747* / 0.734* / 0.934* 42.868* / 0.724* / 0.944* 42.960* / 0.717* / 0.950* 43.079* / 0.707* / 0.957* 43.156* / 0.701* / 0.961*
DiffusionMBIR+2.5D 42.615* / 0.745* / 0.913* 42.834* / 0.727* / 0.935* 42.892* / 0.722* / 0.945* 42.911* / 0.720* / 0.948* 42.931* / 0.719* / 0.952* 42.950* / 0.717* / 0.955*
TPDM 42.691* / 0.739* / 0.906* 42.810* / 0.729* / 0.929* 42.939* / 0.718* / 0.940* 43.051* / 0.709* / 0.946* 43.202* / 0.697* / 0.954* 43.291* / 0.690* / 0.959*
TPDM+2.5D 42.334* / 0.769* / 0.887* 42.458* / 0.759* / 0.912* 42.483* / 0.757* / 0.923* 42.499* / 0.755* / 0.927* 42.528* / 0.753* / 0.932* 42.536* / 0.752* / 0.935*
DDPET-3D (proposed) 52.899 / 0.267 / 0.965 54.937 / 0.215 / 0.977 57.119 / 0.171 / 0.985 58.551 / 0.148 / 0.989 60.916 / 0.103 / 0.993 63.804 / 0.077 / 0.996
@@
Figure 12: Images reconstructed using different methods with and without the proposed 2.5D conditional strategy. Note that all the images were normalized by the total injected activities. Blue arrows point to a possible lesion that is better recovered by DDPET-3D with more conditional slices.
@@
Figure 15: Without the denoised prior, proposed DDPET-3D still produced more consistent reconstructions compared with other methods. Blue arrows point to regions with noticeable inconsistencies between slices in competing methods. Without the denoised prior, DDPET-3D still produced more consistent reconstructions compared with other methods.
@@
Impact of Number of Conditional Slices: We evaluated DDPET-3D with different numbers of neighboring conditional slices. Three variants of the DDPET-3D networks were trained using 9, 21, and 41 neighboring slices as conditional information. These networks are denoted as ”DDPET-3D ($n$ slice)”, where $n$ is the number of conditioned neighboring slices. The results showed that $n=31$ performed the best in most quantitative metrics. Therefore, $n=31$ was used in this paper. Quantitative results are presented in Table IV. To facilitate the comparison, we only conducted the ablation studies on a subset of the patient studies acquired at the Shanghai Ruijin Hospital. As presented in Fig. 11, using more conditioned slices helped the network to recover some subtle details in the images (blue arrows in Fig. 11).
@@
For comparison, we also extended the comparison diffusion models (DiffusionMBIR, TPDM, and DDIM), and re-trained them using neighboring slices as conditional information (denoted as DiffusionMBIR+2.5D, TPDM+2.5D, and DDIM+2.5D). Sample reconstruction results are presented in Fig. 12. Simply adding neighboring slices as conditional information does not necessarily lead to better performance in other baseline models. The corresponding quantitative evaluations are presented in Table V.
@@
Impact of Denoised Prior: We tested the DDPET-3D without the denoised prior during the sampling process to demonstrate the improvement in image quantification using the proposed denoised prior. This network is denoted as ”DDPET-3D (no prior)” in Table IV. As presented in Fig. 13, we can see that diffusion models produced images with inaccurate tracer activities in different organs without the proposed denoised prior (especially in the brain and liver). Note that the images were already normalized by the total injected activities of the entire 3D volume. Also, without the denoised prior, many details in the images were not able to be recovered, as presented in the brain slice in Fig. 13. In addition, we implemented the denoised prior in other competing methods for comparison, so that all the methods have the same starting point in the diffusion process. As shown in Fig. 14, DDPET-3D still outperformed other competing methods, demonstrating the effectiveness of other proposed techniques in DDPET-3D. With the denoised prior, DiffusionMBIR and TPDM produced images with better PSNR and RMSE measurements than the networks without it, but inconsistencies between slices still exist. Lastly, as shown in Fig. 15, without the denoised prior, DDPET-3D still produced more consistent reconstructions compared to other methods. Note that the denoised prior is one of the key contributions proposed in this work.
@@
Impact of Fixing Noise Variables: We analyzed the DDPET-3D without and with fixing the 2 noise variables to demonstrate its effectiveness in producing consistent 3D reconstructions. Specifically, $\epsilon_{0}^{a}$ and $\epsilon_{0}^{b}$ are sampled from ${\mathcal{N}}(0,\textbf{I})$ for each slice in the 3D image volume, instead of fixing them for all slices. This network is denoted as ”DDPET-3D (no fix $\epsilon$)” in Table IV. Although the differences in quantitative evaluations without and with fixing noise variables are small, and not fixing noise even led to better quantitative measurements in certain cases, we noticed significant improvements in visual quality, as presented in Fig. 16. Not fixing noise variables produced images with inconsistent slices and unclear organ boundaries, which are unfavorable in clinical settings.
@@
Impact of Dose Embedding: We tested the DDPET-3D without dose embedding, which is denoted as ”DDPET-3D (no dose)” in Table IV. The experimental results showed that DDPET-3D with dose embedding produced images with better quantitative results.
V Discussion and Conclusion